# Towards Cross-Tokenizer Distillation: the Universal Logit Distillation Loss for LLMs

### v2

@@ 4.1 Evaluation Scenarios
To align with the black-box approach where training models may not be available, we abstain from fine-tuning teacher models. In this way, we enable ULD Loss to operate in an unsupervised environment by generating all answers text with teacher models. For repeatability and fair comparison between experiments, we opted to retain original answers for the test set split. We investigated various scenarios to evaluate the ULD loss performance across different datasets and tasks. These comprised 2 Extractive QA (Ext.), 2 Generative QA (Gen.), and 1 Summary (Sum.) tasks:
SQuAD (Ext.):
The Stanford Question Answering Dataset (SQuAD) Rajpurkar et al. (2016) is a reading comprehension dataset with 87,600 questions generated by crowdworkers from Wikipedia articles. Answers are text portions from the relevant sections of the articles.
QED (Ext.):
The QED Lamm et al. (2020) dataset, expertly annotated, extends from a subset of the Google Natural Questions dataset, comprising 7,640 question-answering pairs with explanations. Our focus is exclusively on extracted answers (spans).
FairytaleQA (Gen.):
The FairytaleQA Dataset Xu et al. (2022), created by educational experts, consists of 10,580 questions from 278 children-friendly stories. Questions may be explicit or implicit.
PubMedQA (Gen.):
The PubMedQA Jin et al. (2019) dataset contains question-answer pairs extracted from medical papers. Questions are based on titles, context on abstracts, and responses on conclusions. Due to the dataset size and context length of our student models, we subset the dataset by selecting the first 50,000 smaller items.
DIALOGSum (Sum.):
DialogSum Chen et al. (2021) is a large-scale dialogue summarization dataset, consisting of 13,460 spoken dialogues with corresponding summaries and topics.
@@ 4.2 Experimental Choices
As far as we know, the only method currently capable of distilling any pair of teacher and student LLM models in a black-box approach is distillation from teacher-generated text seen in Sec. 2.3. Throughout the remainder of this paper, distillation from teacher-generated text will serve as the baseline for evaluating the distillation process using the ULD loss across different teacher-student pairs, tasks, and datasets.
@@
We employed two teacher decoder models, each with 7 billion parameters: LLama 2 7b Chat (LLama) Touvron et al. (2023b) and Mistral 7b Instruct (Mistral) Jiang et al. (2023). These instruct models were chosen for their ability to generate few-shot answers Brown et al. (2020); Wang et al. (2020b) across diverse tasks and their distinct vocabulary set as shown in Figure 1.
@@
We chose student models from various LLM families and architectures with parameters ranging between 160 million to 1 billion: OPT 350m Zhang et al. (2022), Pythia 160m, Pythia 410m, Pythia 1b Biderman et al. (2023), Bloomz 560m Muennighoff et al. (2023) all decoder models and MT0 580m Muennighoff et al. (2023) an encoder-decoder. It’s important to note that models can have been already pre-trained on some datasets such as SQuAD for Bloomz and MT0.
@@
ULD loss distillation and teacher-generated text distillation were processed uniformly. The two teacher models generate answers in inference mode for the five datasets. These answers are then utilized to train student models. During training, student models are trained exclusively to predict answers, in teacher forcing configuration. Logits used for the ULD loss are calculated by applying teacher models to the same data points they generated. Teacher’s weights were frozen, ensuring consistency in teacher-generated sentences during inference and training. Additional parameter details (learning rate, batch size, etc.) can be found in the Appendix Sec. 11.
@@ 4.3 Teacher Performances
Distilling using synthetic teacher-generated answers might restrict student performance on teacher’s ones. To measure distillation efficiency accurately, we report the average native performances across tasks for both teachers Tab. 1 (details in Appendix Sec. 9). We chose a primary metric for each task reflecting associate performances: F1 score for Extractive QA Sokolova et al. (2006), BERTScore for Generative QA Zhang* et al. (2020), and Rouge-Lsum for summary task Lin (2004). Comprehensive evaluation methods and outcomes, encompassing prompts and few-shot examples, are provided in the Appendix Sec. 10.
@@ 5 Empirical Results
Teacher Model Method
SQUAD
QED
FairytaleQA
PubMedQA
DIALOGSum
(Rouge-Lsum)
@@ 5.3 Dataset Size Ablation Study
Figure 4: Dataset size ablation with a LLama/Pythia-410m pair of models trained with ULD loss or teacher-generated text (raw text). The X-axis indicates the % of data used during training while the y-axis represents the test set score for respective datasets.
In this section, we investigate and report in Fig. 3 the influence of the dataset size for models trained with ULD Loss or teacher-generated text. We perform ablations with respectively 25%, 50%, 75%, and 100% of dataset size while keeping training parameters constants. For every ablation ratio, models trained with ULD loss achieved better performance than models trained on teacher-generated text. Specifically, with 50% of a dataset, ULD loss models overpass the performance of teacher-generated text models trained with full dataset.
@@ 5.4 Training Regularization
To understand the impact of the ULD loss during training we decide to compute the validation ULD and Cross-entropy loss values for two pairs of teacher/student on the SQuAD dataset every 200 steps during 5 epochs. We report the curves formed by this point in Fig. 5.
Figure 5: Evolution of validation ULD and Cross-entropy loss curves during training on SQuAD dataset for a LLama/Pythia-410m and LLama/Bloomz-560m Teacher/Student pair of model. For teacher-generated text models (raw text), the ULD loss was only computed during validation and did not impact the training process.
It appears that using the ULD loss contributes to stabilizing the distillation process over training and mitigates overfitting issues, enabling the model to train more effectively across multiple epochs. It’s worth noting that incorporating the ULD loss during training stabilizes both ULD and Cross-entropy loss.
@@
Higher Cross-entropy loss validation values observed with ULD loss are a direct result of the training process. ULD loss trained model to predict soft labels from the teacher model instead of the one-hot vectors used for computing the Cross-entropy loss.
@@ 6 Distillation of Decoder Teacher to Encoder-Decoder Student
As shown in Sec. 5, ULD loss effectively transfers knowledge from any pair of teacher/student decoders. Moreover, by leveraging solely on logit information and adopting a black-box approach, ULD loss should be able to extend its versatility and improve cross-architecture distillation. To validate this, we distill a teacher/student pair LLama/MT0-580m and focus our experimentation on PubMedQA, DIALOGSum, and QED to avoid any data seen during the pre-training of MT0 with the xP3 dataset Muennighoff et al. (2023); Sanh et al. (2022).
@@
Table 3: Distillation of a LLama teacher (decoder) to an MT0-580m (encoder-decoder) with ULD Loss and teacher-generated text on three data sets and their primary metric associate.
@@ 7 Conclusions
In this work, we introduce the Universal Logit Distillation (ULD) loss, a novel method for distilling any decoder teacher model into any student model for LLM generative tasks. ULD achieves better overall results and matches the performance of teacher-generated text distillation with only half of the training dataset or student size, while effectively preventing overfitting. Our extensive experiments validate the efficacy of the ULD loss across diverse tasks,- datasets, and architectures, demonstrating its superiority over standard teacher-generated text distillation methods.
Limitations
One unexplored avenue related to our work is the study of the ULD loss in non-English languages, including ones with vastly different token representations (Chinese, Korean, etc). However, we believe that our findings can still be applied to other languages, especially if teacher-student model pairs are pre-trained in the same language or achieve multilingual capacity. Additionally, in line with previous work, we evaluate task performance with standard reference-based metrics (ROUGE, F1, BERTScore) which can be limited in their assessment of what constitutes a correct generative model prediction Faysse et al. (2023). An extension of the work could be to go beyond mono-task distillation, and assess the transferability of generalist assistant abilities from larger to smaller models.
Ethics Statement

### v3

@@ 4.1 Evaluation Scenarios
We avoid fine-tuning teacher models to ensure alignment with the black-box approach as training teacher models may be unavailable. In this way, we enable ULD Loss to operate in an unsupervised environment by generating all answers text with teacher models. For repeatability and fair comparison between experiments, we opted to retain original answers for the test set split. We investigated various scenarios to evaluate the ULD loss performance across different datasets and tasks. These comprised 2 Extractive QA (Ext.), 2 Generative QA (Gen.), and 1 Summary (Sum.) tasks:
    SQuAD (Ext.): The Stanford Question Answering Dataset (SQuAD) (Rajpurkar et al., 2016) is a reading comprehension dataset with 87,600 questions generated by crowdworkers from Wikipedia articles. Answers are text portions from the relevant sections of the articles.
    QED (Ext.): The QED (Lamm et al., 2020) dataset, expertly annotated, extends from a subset of the Google Natural Questions dataset, comprising 7,640 question-answering pairs with explanations. Our focus is exclusively on extracted answers (spans).
    FairytaleQA (Gen.): The FairytaleQA Dataset (Xu et al., 2022), created by educational experts, consists of 10,580 questions from 278 children-friendly stories. Questions may be explicit or implicit.
    PubMedQA (Gen.): The PubMedQA (Jin et al., 2019) dataset contains question-answer pairs extracted from medical papers. Questions are based on titles, context on abstracts, and responses on conclusions. Due to the dataset size and context length of our student models, we subset the dataset by selecting the first 50,000 smaller items.
    DIALOGSum (Sum.): DialogSum (Chen et al., 2021) is a large-scale dialogue summarization dataset, consisting of 13,460 spoken dialogues with corresponding summaries and topics.
@@ 4.2 Experimental Choices
As far as we know, the only method currently capable of distilling any pair of teacher and student LLM models in a black-box approach is distillation from teacher-generated text seen in Sec. 2.3. Throughout the remainder of this paper, distillation from teacher-generated texts will serve as the baseline for evaluating the distillation process using the ULD loss across different teacher-student pairs, tasks, and datasets. Additional experimentation with KL Div methods between models from similar families can be found in Appendix Sec. B.2
@@
We employed two teacher decoder models, each with 7 billion parameters: LLama 2 7b Chat (LLama) by Touvron et al. (2023b) and Mistral 7b Instruct (Mistral) by Jiang et al. (2023). These instruct models were chosen for their ability to generate few-shot answers (Brown et al., 2020; Wang et al., 2020b) across diverse tasks and their distinct vocabulary set as shown in Figure 1.
@@
We chose student models from various LLM families and architectures with parameters ranging between 160 million to 1 billion: OPT 350m (Zhang et al., 2022), Pythia 160m, Pythia 410m, Pythia 1b (Biderman et al., 2023), Bloomz 560m (Muennighoff et al., 2023) all decoder models and MT0 580m (Muennighoff et al., 2023) an encoder-decoder. It’s important to note that models can have been already pre-trained on some datasets such as SQuAD for Bloomz and MT0.
@@
ULD loss distillation and teacher-generated text distillation were processed uniformly. The two teacher models generate answers in inference mode for the five datasets. These answers are then utilized to train student models. During training, student models are trained exclusively to predict answers, in teacher forcing configuration. Logits used for the ULD loss are calculated by applying teacher models to the same data points they generated. Teacher’s weights were frozen, ensuring consistency in teacher-generated sentences during inference and training. Additional parameter details (learning rate, batch size, etc.) can be found in the Appendix App. G.
@@ 4.3 Teacher Performances
Distilling using synthetic teacher-generated answers might restrict student performance on teacher’s ones. To measure distillation efficiency accurately, we report the average native performances across tasks for both teachers Tab. 1 (details in Appendix App. D). We chose a primary metric for each task reflecting associate performances: F1 score for Extractive QA (Sokolova et al., 2006), BERTScore for Generative QA (Zhang* et al., 2020), and Rouge-Lsum for summary task (Lin, 2004). Comprehensive evaluation methods and outcomes, encompassing prompts and few-shot examples, are provided in the Appendix App. E.
@@ 5 Empirical Results
Teacher Model Method SQUAD QED FairytaleQA PubMedQA DIALOGSum
(F1) (F1) (BERTScore) (BERTScore) (Rouge-Lsum)
@@ 5.3 Dataset Size Ablation Study
In this section, we investigate and report in Fig. 4 the influence of the dataset size for models trained with ULD Loss or teacher-generated text. We perform ablations with respectively 25%, 50%, 75%, and 100% of dataset size while keeping training parameters constants. Training for each ratio was conducted five times to establish a range of performance. For every ablation ratio, models trained with ULD loss achieved better performance than models trained on teacher-generated text. Specifically, with 50% of a dataset, ULD loss models overpass the performance of teacher-generated text models trained with full dataset.
Figure 4: Dataset size ablation with a LLama/Pythia-410m pair trained with ULD loss or teacher-generated text. The X-axis indicates the % of data used during training while the y-axis represents the test set score. Minimum and maximum values are represented by the error bars in the plot while the mean is represented by points.
@@ 5.4 Training Regularization
To understand the impact of the ULD loss during training we decide to compute the validation ULD and Cross-entropy loss values for two pairs of teacher/student on the SQuAD dataset every 200 steps during 5 epochs. We report the curves formed by this point in Fig. 5. It appears that using the ULD loss contributes to stabilizing the distillation process over training and mitigates overfitting issues, enabling the model to train more effectively across multiple epochs. It’s worth noting that incorporating the ULD loss during training stabilizes both ULD and Cross-entropy loss.
Figure 5: Evolution of validation ULD and Cross-entropy loss curves during training on SQuAD dataset for a LLama/Pythia-410m and LLama/Bloomz-560m Teacher/Student pair of model. For teacher-generated text models (raw text), the ULD loss was only computed during validation and did not impact the training.
@@
Even though the cross entropy is lower with raw text data than with ULD Loss, the raw data does not lead to better results. In fact, during training, models distilled with ULD Loss are trained to reproduce the predictions of the teacher model. Consequently, they are not trained to minimize cross-entropy and have maximum confidence in each prediction.
For this reason, a student model trained to reproduce the confidence of a teacher model of 0.92, compared to a model trained only on raw text and trained to predict 1, will have a lower probability on the golden token and consequently a higher Cross Entropy Loss. However, this model will be better calibrated to its true confidence for the sentence.
@@ 6 Distillation of Decoder Teacher to Encoder-Decoder Student
As shown in Sec. 5, ULD loss effectively transfers knowledge from any pair of teacher/student decoders. Moreover, by leveraging solely on logit information and adopting a black-box approach, ULD loss should be able to extend its versatility and improve cross-architecture distillation. To validate this, we distill a teacher/student pair LLama/MT0-580m and focus our experimentation on PubMedQA, DIALOGSum, and QED to avoid any data seen during the pre-training of MT0 with the xP3 dataset (Muennighoff et al., 2023; Sanh et al., 2022).
Table 3: Distillation of a LLama teacher (decoder) to an MT0-580m (encoder-decoder) with ULD Loss and teacher-generated text on three data sets.
@@ 7 Conclusions
In this work, we introduce the Universal Logit Distillation (ULD) loss, a novel approach for distilling any decoder teacher model into any student model for LLM generative tasks, utilizing a new closed form of the Wasserstein distance never seen before for distillation. ULD achieves superior overall results and matches the performance of traditional teacher-generated text distillation with only half the training dataset or student model size, while effectively preventing overfitting. Our comprehensive experiments demonstrate the efficacy of the ULD loss across a variety of tasks, datasets, and model architectures, showcasing its advantages over standard teacher-generated text distillation methods.
Broader Impact Statement